Using these data from the comparative balance shee
Answer #1
2014 2013 Amount Percentage
Accounts Receiveble     4,60,000     4,00,000 460,000-400,000 = 60,000 (60,000/400,000)*100 = 15%
Inventory     7,80,000     6,50,000 780,000-650,000 = 130,000 (130,000/650,000)*100 = 20%
Total Assets 31,64,000 28,00,000 3,164,000-2,800,000 = 364,000 (364,000/2,800,000)*100 = 13%